OpenStudy (anonymous):

The formula for the surface area of a rectangular prism with a square base is SA = 2s^2 + 4sh. What is the surface area of this rectangular prism if s = 2 and h = 4?

OpenStudy (anonymous):

\[\large S(s,h)=2s^2+4sh \implies S(2,4)=2\times 2^2+4\times(2\times 4)=40\]

OpenStudy (anonymous):

thanks !!